OpenStudy (anonymous):

brad is going on a vacation w/his family. he will be going thru six states on his trip, but traveling only 900 miles. the vacation will last twenty three days. while brad is happy to be missing school, his teacher asked him to write in his journal one half page each day of his trip. how many pages will brad have written when he gets back home? EXPLAIN

OpenStudy (anonymous):

23*1.5=34.5 pages

OpenStudy (dumbcow):

1/2 page each day for 23 days --> 23*1/2 = 11.5 pages

OpenStudy (anonymous):

half page is 0.5 one half is 1.5

OpenStudy (anonymous):

if its 0.5 than ur ans is right 11.5

OpenStudy (anonymous):

one half = 0.5 23*0.5 = 11.5 pages

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

kk thanks:D